I throw a rock horizontally off of a cliff at 25 m/s. If the cliff is 120 m tall, how far away will the rock land?
1
Expert's answer
2020-11-05T10:44:10-0500
"x=vt=v\sqrt{\frac{2h}{g}}\\x=25\sqrt{\frac{2(120)}{9.8}}=124\ m"